# Preference Distillation for Personalized Generative Recommendation

### v1

@@ 4.5 Training and Generative Recommendation
Using the task-alternated training strategy proposed by Li et al. (2023), we alternate the task trained for each batch. This allows us to save time during training because the samples in the batch have the same data format and will therefore be a similar length to one another. Importantly, task-alternated training did not show any degradation in generative recommendation performance versus traditional random batching strategies Li et al. (2023). We combine user-personalized prompts, task-specific prompts, and discrete prompts as input during training. We also add whole-word embeddings to help the model identify which tokens belong to a particular ID. We show the whole mode architecture in Figure 2.
@@
$\mathcal{L}_{\Theta}=\frac{1}{|\mathcal{D}|}\sum_{(X,Y)\in\mathcal{D}}\frac{1}{|Y|}\sum_{t=1}^{|Y|}-\log p(y_{t}|Y_{<t},X)$ (4)
@@
In addtion, the model must output a sequence of text that forms an item ID (or a natural language text in the case of explanation generation). For this task, we choose beam search, which is commonly used for its effectiveness in sequence-to-sequence generation. Suppose we set the number of beams to $b$. At each time step, are $b$ candidate sequences. In the next step, any word in the vocabulary $V$ are appended to the end of the candidate sequences, resulting in $b\times V$ combinations. We can then select $b$ sequences that have the maximum log-likelihood. The LLM can continue this process until the candidate sequences reach a predefined maximum length. For sequential recommendation and top-$N$ recommendation, the $b$ candidate sequences form the recommendation list. For explanation, we select the sequence with the largest log-likelihood from the candidates.
5 Experimental Setup
5.1 Datasets
For experimentation, we use the Sports, Beauty, and Toys & Games Dataset from Amazon  [^1] https://jmcauley.ucsd.edu/data/amazon/. We reuse the preprocessing steps and the 8:1:1 train/validation/test splits created by Geng et al. (2022). Each sample consists of a user ID, an item ID, a text review, a rating, and a time stamp. We provide the dataset statistics in Table 1. For explanation generation, the Sentires toolkit Zhang et al. (2014) was used to extract ground-truth explanations for a given (user, item) pair. For sequential recommendation, the last item in a user’s history is part of the test set, the penultimate item is part of the validation set, and all other items are included in the train set. To prevent data leakage, the same training split used for sequential recommendation is used for top-n recommendation. For reproducibility, we provide all code used in the supplementary materials section of our submission and will release the code publicly upon publication.
@@ 5.1 Datasets
Table 1: Statistics of the datasets.
5.2 Baselines
For sequential recommendation, we compare our model with the following baselines: Caser Tang and Wang (2018), HGN Ma et al. (2019), GRU4Rec Jannach and Ludewig (2017), BERT4Rec Sun et al. (2019), FDSA Zhang et al. (2019), SASRec Kang and McAuley (2018), P5 Geng et al. (2022), VIP5 Geng et al. (2023), and POD Li et al. (2023).
For top-n recommendation, we compare our method with: MF Koren et al. (2009), MLP Cheng et al. (2016), P5 Geng et al. (2022), VIP5 Geng et al. (2023), and POD Li et al. (2023).
Finally, for explanation generation, we use the following baselines: Att2Seq Dong et al. (2017):, NRT Li et al. (2020), PETER Li et al. (2021), and POD Li et al. (2023). For a fair comparison, we exclude P5 and VIP5 because we are focused exclusively on models that only use user and item IDs rather than additional user or item metadata.
Table 2: Performance comparison on sequential recommendation.
@@ 5.2 Baselines
PeaPOD 0.0505 0.0400 0.0611 0.0432 0.0588 0.0445 0.0738 0.0493 0.0692 0.0583 0.0787 0.0609
Table 3: Performance comparison on top-n recommendation.
@@
PeaPOD 0.1212 0.2741 0.2007 0.3616 0.2286 0.1097 0.2441 0.1780 0.3260 0.2039 0.0728 0.1618 0.1190 0.2315 0.1411
Table 4: Performance comparison on explanation generation.
@@
PeaPOD 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
5.3 Evaluation Metrics
For sequential and top-n recommendation, we evaluate models using two commonly used metrics, normalized Discounted Cumulative Gain (nDCG) and Hit Ratio (HR). nDCG is a metric used to evaluate the quality of ranked lists by considering the position of relevant items, while HR measures the fraction of relevant items that appear in the top-n positions of a ranked list. In particular, we measure HR@{5,10} and nDCG@{5,10} for sequential recommendation and HR@{1,5,10} and nDCG@{5,10} for top-n recommendation. For explanation generation, we evaluate models using BLEU Papineni et al. (2002) and ROUGE Lin and Hovy (2003). BLEU evaluates machine translation quality using n-gram precision against reference translations, while ROUGE assesses text summarization and translation by comparing n-gram overlaps with reference texts. We report the F1 scores for BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L. For all metrics, a greater score indicates better performance. In all tables, the best performing model for each metric is in bold and the second best model is underlined.
5.4 Implementation Details
To ensure a fair comparison, we compare PeaPOD with POD Li et al. (2023), P5 Geng et al. (2022), and VIP5 Geng et al. (2023) using T5-small (60.5 M parameters) Raffel et al. (2020) as the pre-trained LLM. On average, each experiment took 6 hours to run. Since the main focus of our study is not on formulating discrete prompts, we reuse the same discrete prompts as Geng et al. (2022). We adapt task-alternated training, as proposed by Li et al. (2023), to speed up training time by alternating each task per batch. Notably, both VIP5 and P5 both mention that fine-tuning with the T5-small pre-trained checkpoint often outperforms fine-tuning with T5-base.
For all models, we train the model on an Nvidia RTX 3090 for 30 epochs, a batch size of 64, the AdamW optimizer Kingma and Ba (2015), and a learning rate of 5e-3. We save the best performing model on the validation set. The length of the task-specific prompts and prompt components are both set to 3. We set the number of prompt components to 20. For our hyperparameter search, we tested prompt lengths of {3, 5, 8, 10} and the number of prompts of {10, 20, 30, 100}.To obtain the user embedding, we train a probabalistic matrix factorization model Mnih and Salakhutdinov (2007), a variant of matrix factorization, for 100 iterations with an embedding size of 512, learning rate of 1e-3, and a lambda regularization of 1e-3. During inference time, the number of beams is set to 20 and the inference batch size is set to 32 for all tasks.
6 Results
6.1 Sequential Recommendation
In Table 2, we compare our models with our selected baseline models. Our results show that PeaPOD outperforms all baselines in the Sports and Beauty datasets and is either the first or second best model for all metrics in the Toys dataset. Notably, the only difference between our method and POD, which is the second best performing model, is the additional personalized-user prompt that is prepended to the input. Interestingly, even though the user embedding passed to the attention mechanism is trained on a top-n recommender model, the additional information still provides a performance boost to sequential recommendation. This performance gain may be attributed to the way the data was split. In particular, the sequential and top-n recommendation contains the same user-item interactions, meaning that information about the users can be shared to improve performance in both tasks.
6.2 Top-n Recommendation
In Table 3, we report the performance of the top-n recommendation task. We note that PeaPOD performs the best in all metrics for all datasets, with a significant improvement in performance. We note that matrix factorization, which is used to generate the user embedding, is trained for top-n recommendation. Thus, the boost in performance may be more significant when compared to the other two tasks because of the user embedding selected as the query function. Nevertheless, we also find that PeaPOD outperforms matrix factorization itself, meaning that combining traditional collaborative filtering methods with LLMs leads to increased performance.
@@ 6.2 Top-n Recommendation
Figure 3: Comparison between PeaPOD and an ablated version (PeaPOD-X) trained without the task-specific prompts on our three tasks.
6.3 Explanation Generation
In Table 4, we report the performance of all models on the explanation generation task. Interestingly PeaPOD performs particularly well on both Sports and Toys datasets, but performs significantly worse in Beauty when compared to the top baseline models. Nevertheless, PeaPOD ranks first or second best on six out of the twelve evaluations, meaning that it is still competitive in explanation generation.
Upon manual inspection across all models, we notice that there are a significant amount of duplicates across explanations, meaning that the same explanations are generated regardless of the (user, item) pair. Thus, we posit that limiting the explanation generation to exclusively user and item IDs hurts performance. Similar to how performance for sequential and top-n recommendation was improved by providing collaborative user embeddings, the performance of item explanations can be improved by providing additional item metadata to the model. Previous research Geng et al. (2023); Tan et al. (2024); Rajput et al. (2023) has already explored how discrete item IDs can be transformed to provide more informative tokens to the model. However, we leave the additional experiments of combining enhanced item IDs with PeaPOD to future work.
6.4 Training without Task-Specific Prompts
@@
To understand the effectiveness of user-personalized prompts versus globally shared prompts, we fine-tune the model with the same architecture and hyperparameters, but omit the task-specific prompts. We refer to the ablated model as ‘PeaPOD-X’.
As shown in Figure 3, we find that in all three datasets, the performance of the sequential recommendation marginally decreases, but the top-n recommendation performance improves. For explanation generation, the original model performs better for the Beauty dataset, but worse for Sports and Toys. We attribute the boost in top-n recommendation performance to the user embedding used for the query function. In particular, matrix factorization is a top-n recommendation model, meaning that the model is most likely to excel in this task. However, the lack of a task-specific prompt for sequential recommendation hinders the performance on this task. Nevertheless, we observe for the sequential recommendation task, the ablated model is still better than solely using the discrete prompt and has similar performance to POD. Thus, we show the effectiveness of a user-personalized prompt that utilizes preferences shared between similar users to improve recommendation performance. We report all metrics for this experiment in Appendix B.
7 Conclusion
In this paper, we propose PeaPOD, a novel recommender model that dynamically weights a set of decomposed prompt components to distill user preferences in a compositional manner. We maintain a many-to-many mapping between user preferences and continuous prompts to enable knowledge sharing among users with similar preferences. We utilize an attention-based mechanism to weight prompt components based on the user’s interaction history, thus creating a user-personalized prompt that distills both the user’s and similar users’ preferences into a learnable soft prompt. In our experiments, we show that our user-personalized prompt is effective in increasing the performance for three common recommendation tasks. Our model shows the effectiveness of preference distillation to improve personalized generative recommendation.
8 Limitations and Ethical Concerns
One limitation to our approach is that we limit our study to exclusively use user and item IDs. Thus, we do not incorporate any additional metadata such as user profiles or item descriptions. We find that this may hinder performance on explanation generation, which relies on having a strong understanding of both the user and item to provide a personalized response. We predict that providing additional metadata to the model can improve performance, particularly if it is given in natural language text as the LLM is already pre-trained on such data. We plan to extend PeaPOD to utilize this information in future work.
In addition, our model takes a two-step approach to generate user-personalized prompts. Namely, the user embedding is generated using matrix factorization, which is then passed to the model as the query function. However, the models are not connected end-to-end, which means that the user embedding cannot be continuously updated during training. For future work, we plan to combine these two models into a single end-to-end framework.
Another concern is the issue of fairness in generative recommender systems. For example, prior research has shown that LLM-based recommenders skew towards popular items, which causes a disparity in recommendations and can lead to a bias-amplification loop Yoon et al. (2024). In addition, past research has investigated the issue of fairness in the LLM itself Salutari et al. (2023). Thus, careful consideration must be made when deploying a generative recommender model to mitigate unfair recommendations.
Appendix A Baselines
For each of our three tasks, we compare our model with the following baselines.
A.1 Sequential Recommendation
Sequential recommendation predicts a user’s next preferred items based on their previous interactions, which are given in historical order.
  • Caser: ConvolutionAl Sequence Embedding Recommendation Tang and Wang (2018) uses a convolutional neural network to learn the sequential pattern of the user interaction history.
  • HGN Ma et al. (2019): Hierarchical Gating Network uses two gating modules to model a user’s short and long term interests.
  • GRU4Rec Jannach and Ludewig (2017): Gated Recurrent Unit (GRU) for Recommendation uses a GRU to process a user’s recommender session.
  • BERT4Rec Sun et al. (2019): Fine-tunes a bidirectional encoder representations from transformers model for sequential recommendation.
  • FDSA Zhang et al. (2019): Feature-level Deeper Self-Attention network. Integrates item features as feature sequences, which are then combined with item sequences to generate recommendations.
  • SASRec Kang and McAuley (2018): Self-Attentive based Sequential Recommendation model. Uses self-attention mechanisms for short-term semantics and Recurrent Neural Networks for long-term semantics.
  • $\mathbf{S^{3}}$-Rec Zhou et al. (2020): Self-Supervised learning for Sequential Recommendation. Uses four self-supervised learning objectives to capture the correlations between items in sequential data.
[…]
  • VIP5 Geng et al. (2023): An extension of P5 that trains on multimodal data.
  • POD Li et al. (2023): An extension of P5 that uses a task-specific, distilled prompt as part of the input.
A.2 Top-n Recommendation
Top-n recommendation predicts a list of the top-n items based on the user’s interaction history.
  • MF Koren et al. (2009): Matrix Factorization is a collaborative filtering technique that decomposes a user-item interaction matrix into latent user and item embeddings.
  • MLP Cheng et al. (2016): Multi-Layer Perceptron is a neural network trained to learn captures non-linear interactions between users and items.
  • P5 Geng et al. (2022): See definition in Section A.1.
  • VIP5 Geng et al. (2023): See definition in Section A.1.
  • POD Li et al. (2023): See definition in Section A.1.
A.3 Explanation Generation
Given a user ID and item ID, the model must return a natural language explanation describing why the user might like/dislike the target item. Notably, we exclude P5 and VIP5 because we are exclusively focused on models that exclusively use only user and item IDs rather than item titles or other item metadata.
  • Att2Seq Dong et al. (2017): an attention based model that learns to generate product reviews from item attributes.
  • NRT Li et al. (2020): Neural Ratings and Tips generation simultaneously returns a rating and a tip based on a neural network.
  • PETER Li et al. (2021): a transformer based model that performs rating prediction and explanation generation.
Appendix B Training without Task-Specific Prompts
We report the original model including task-specific prompts (denoted as PeaPOD) versus training the model without them (denoted as PeaPOD-X). All inputs and hyperparameters are otherwise the same.

### v2

@@ 4.4. Training and Generative Recommendation
Using the task-alternated training strategy proposed by Li et al. (2023), we alternate the task trained for each batch. This allows us to save time during training because the samples in the batch have the same data format and will therefore be a similar length to one another. Importantly, task-alternated training did not show any degradation in generative recommendation performance versus traditional random batching strategies (Li et al., 2023). We combine collaborative user prompts, task-specific prompts, and discrete prompts as input during training. We also add whole-word embeddings to help the model identify which tokens belong to a particular ID (Geng et al., 2022). We show the whole model architecture in Figure 2.
@@
(10) $\mathcal{L}_{\Theta}=\frac{1}{|\mathcal{D}|}\sum_{(X,Y)\in\mathcal{D}}\frac{1}{|Y|}\sum_{t=1}^{|Y|}-\log p(y_{t}|Y_{<t},X)$
@@
In addition, the model must output a sequence of text that forms an item ID (or natural language text in the case of explanation generation). For this task, we choose beam search, which is commonly used for its effectiveness in sequence-to-sequence generation. Suppose we set the number of beams to $b$. At each time step, are $b$ candidate sequences. In the next step, any word in the vocabulary $\boldsymbol{V}$ are appended to the end of the candidate sequences, resulting in $b\times\boldsymbol{V}$ combinations. We can then select $b$ sequences that have the maximum log-likelihood. The LLM can continue this process until the candidate sequences reach a predefined maximum length. For sequential recommendation and top-n recommendation, the $b$ candidate sequences form the recommendation list. For explanation, we select the sequence with the largest log-likelihood from the candidates.
5. Experimental Setup
The following research questions will guide the remainder of the paper:
  • RQ1: Does our proposed PeaPOD model outperform existing state-of-the-art models on different tasks?
  • RQ2: Does using a different method to obtain user embeddings significantly change the recommendation performance?
  • RQ3: What is the effectiveness of adding a collaborative user prompt as input to the model?
  • RQ4: How does model size impact performance?
  • RQ5: How useful is multi-head attention vs. alternative implementations such as single-head attention or multi-layer perceptron?
5.1. Datasets
To answer the research questions, we use the Sports, Beauty, and Toys & Games Dataset from Amazon (Ni et al., 2019). We reuse the preprocessing steps and the 8:1:1 train/validation/test splits created by Geng et al. (2022). Each sample consists of a user ID, an item ID, a text review, a rating, and a time stamp. We provide the dataset statistics in Table 2. For explanation generation, the Sentires toolkit (Zhang et al., 2014) was used to extract ground-truth explanations for a given (user, item) pair. For sequential recommendation, the last item in a user’s history is part of the test set, the penultimate item is part of the validation set, and all other items are included in the train set. To prevent data leakage, the same training split used for sequential recommendation is used for top-n recommendation.
@@ 5.1. Datasets
Table 2. Statistics of the datasets.
5.2. Baselines
We report the following baselines for each of the three recommendation tasks:
  • For sequential recommendation, we compare with the following baselines: Caser (Tang and Wang, 2018), HGN (Ma et al., 2019), GRU4Rec (Jannach and Ludewig, 2017),
    BERT4Rec (Sun et al., 2019), FDSA (Zhang et al., 2019), SASRec (Kang and McAuley, 2018), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023), POD (Li et al., 2023), IDGenRec (Tan et al., 2024a), and RDRec (Wang et al., 2024).
  • For top-n recommendation, we compare our method with: MF (Koren et al., 2009), MLP (Cheng et al., 2016), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023), POD (Li et al., 2023), and RDRec (Wang et al., 2024).
  • For explanation generation, we use the following baselines: Att2Seq (Dong et al., 2017):, NRT (Li et al., 2020), PETER (Li et al., 2021), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023) and POD (Li et al., 2023).
5.3. Evaluation Metrics
For sequential and top-n recommendation, we evaluate models using two commonly used metrics, normalized Discounted Cumulative Gain (nDCG) and Hit Ratio (HR). nDCG is a metric used to evaluate the quality of ranked lists by considering the position of relevant items, while HR measures the fraction of relevant items that appear in the top-n positions of a ranked list. In particular, we measure HR@{5,10} and nDCG@{5,10} for sequential recommendation and HR@{1,5,10} and nDCG@{5,10} for top-n recommendation. For explanation generation, we evaluate models using BLEU (Papineni et al., 2002) and ROUGE (Lin and Hovy, 2003). BLEU evaluates machine translation quality using n-gram precision against reference translations, while ROUGE assesses text summarization and translation by comparing n-gram overlaps with reference texts. We report the F1 scores for BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L. For all metrics, a greater score indicates better performance. In all tables, the best performing model for each metric is in bold and the second best model is underlined.
5.4. Implementation Details
To ensure a fair comparison, we compare PeaPOD with POD (Li et al., 2023), P5 (Geng et al., 2022), and VIP5 (Geng et al., 2023) using T5-small (60.5 M parameters) (Raffel et al., 2020) as the base LLM. On average, each experiment took 6 hours to run. Since the main focus of our study is not on formulating discrete prompts, we reuse the same discrete prompts as POD (Li et al., 2023). We adapt task-alternated training, as proposed by Li et al. (2023), to speed up training time by alternating each task per batch.
To initialize user embeddings for PeaPOD, we train a probabilistic matrix factorization model (Mnih and Salakhutdinov, 2007) for top-n recommendation. We set the hyperparameters to 100 iterations with a dimension size of 512, learning rate of 1e-3, and a lambda regularization of 1e-3. We will refer to this model as PeaPOD-PMF. We train PeaPOD-PMF on an Nvidia RTX 3090 for 30 epochs. We utilize the AdamW optimizer (Kingma and Ba, 2015), with a batch size of 64, learning rate of 5e-3, and length of the task-specific prompt set to 3. We save the best performing model on the validation set.
For the collaborative user prompt, we perform a hyperparameter search, testing prompt lengths of {3, 5, 8, 10}, number of heads of {1, 2, 4, 8, 16} and the number of similar users of {5, 10, 20, 30}. For Sports and Beauty, the optimal hyperparameters we found was 20 similar users and the length of the collaborative user prompt $|\boldsymbol{p}_{mha}|=3$. For Toys, the optimal hyperparameters we found was 10 similar users and $|\boldsymbol{p}_{mha}|=3$. For all dataset, $H=4$. During inference time, the number of beams is set to 20 and the inference batch size is set to 32 for all tasks. We include our implementation at: https://anonymous.4open.science/r/peapod-3FF7/.
6. Results
6.1. Recommendation Performance (RQ1)
6.1.1. Sequential Recommendation
In Table 3, we compare our models with the baseline models on the sequential recommendation task. Our results show that PeaPOD-PMF outperforms all baselines in the three datasets. In addition, the superior performance of our model shows that user embeddings generated from one recommendation task can act as a strong signal for other recommender tasks. Although the set of user embeddings is generated from PMF, a top-n recommendation model, PeaPOD-PMF outperforms traditional sequential recommendation models. Thus, PeaPOD can effectively utilize personalized information from PMF to perform recommendation tasks that PMF is not explicitly trained for.
Table 3. Performance comparison on sequential recommendation.
@@ 6.1.1. Sequential Recommendation
RDRec 0.0505 0.0408 0.0596 0.0433 0.0601 0.0461 0.0743 0.0504 0.0723 0.0593 0.0802 0.0605
PeaPOD-PMF 0.0534 0.0424 0.0638 0.0456 0.0718 0.0527 0.0915 0.0591 0.0865 0.0708 0.1018 0.0750
Table 4. Performance comparison on top-n recommendation.
@@
RDRec 0.1285 0.2747 0.2033 0.3683 0.2326 0.1203 0.2572 0.1902 0.3380 0.2160 0.0660 0.1655 0.1171 0.2375 0.1398
PeaPOD-PMF 0.1252 0.2708 0.2006 0.3596 0.2287 0.1292 0.2628 0.1978 0.3423 0.2228 0.0728 0.2126 0.1581 0.2872 0.1817
Table 5. Performance comparison on explanation generation.
@@
PeaPOD-PMF 1.0313 13.7253 1.8260 10.7078 1.0468 12.9759 1.8853 10.2740 2.5145 13.8132 4.0357 11.5514
6.1.2. Top-n Recommendation
In Table 4, we report the performance of our method and the baselines on the top-n recommendation task. We found that PeaPOD-PMF outperformed existing baselines on the Beauty and Toys datasets and placed second in Sports. Although PeaPOD-PMF was initialized on a variant of matrix factorization, PeaPOD-PMF showed significant gains over the matrix factorization baseline. Thus, our attention-based compositional strategy is able to further refine the collaborative information provided by the initial user embedding. We explore the effects of the selected method to initialize user embeddings in Section 6.2.
6.1.3. Explanation Generation
In Table 5, we report the performance of all models on the explanation generation task. We observe that PeaPOD-PMF performs competitively in Toys and Sports. In particular, PeaPOD-PMF achieved the best performance in BLEU-4 and ROUGE-2 in Toys. However, PeaPOD-PMF performs worse in Beauty. Importantly, the top-two models, P5 and VIP5, contained additional item metadata in the input that the other models did not have. In particular, P5 provides the item title in the prompt and VIP5 includes both the item title and the item photo. Thus, we posit that including item metadata in the input greatly contributes to the improvement of performance for explanation generation. Although the scope of our study focuses on scenarios where only user and item IDs are available to the model, PeaPOD can be easily extended to include these additional metadata as input. Integrating these complementary methods can lead to better overall performance, making it a compelling avenue for future research.
6.2. Sensitivity to User Embeddings (RQ2)
We train an ablated variant of PeaPOD, named PeaPOD-BiVAE, to understand the effect that initial user embeddings have on the PeaPOD architecture. We initialize the user embeddings of PeaPOD-BiVAE by training a Bilateral Variational Autoencoder (BiVAE) (Truong et al., 2021) for top-n recommendation. We train BiVAE for 100 epochs with an embedding size of 512, learning rate of 1e-3, and $\lambda$ of 1e-3. We report the results of both PeaPOD models in Figure 3.
Overall, we find that the method of initializing user embeddings affects how well the model performs on the different tasks. PeaPOD-BiVAE performed better on sequential recommendation for Sports, but PeaPOD-PMF performed better on all other metrics. Thus, the method used to obtain the initial user embeddings plays an important role in the overall recommendation performance.
It is important to note that the user embedding does not need to be generated from a traditional recommender system. The primary reason we chose to initialize user embeddings using collaborative filtering was that it provides key information about user-to-user interactions that the attention mechanism can further refine. However, any method that provides a strong signal about personalized and collaborative preferences can be used as input to PeaPOD. For future work, we plan to optimize the user embedding initialization and collaborative user prompt generation steps into a unified, end-to-end architecture so that the two components can be trained at the same time rather than splitting them into two distinct steps.
Figure 3. Comparison between PeaPOD-PMF and PeaPOD-BiVAE.
@@ 6.2. Sensitivity to User Embeddings (RQ2)
Figure 4. Comparison between PeaPOD-PMF and an ablated version, PeaPOD (no task), trained without task-specific prompts.
6.3. Effects of Collaborative User Prompt (RQ3)
@@
To understand the effectiveness of collaborative user prompts versus globally shared prompts, we fine-tune PeaPOD-PMF with the same architecture and hyperparameters, but omit the task-specific prompts. We refer to the ablated model as PeaPOD-PMF (no task). We report our findings in Figure 4.
We find that for sequential and top-n recommendation, PeaPOD-PMF performs better on Beauty and Toys, whereas PeaPOD-PMF (no task) performs better in Sports. For explanation generation, the models perform similar to each other, with a slight edge for PeaPOD-PMF in Beauty and PeaPOD-PMF (no task) in Toys. Importantly, we find that PeaPOD-PMF (no task) shows better performance overall versus POD in sequential recommendation and top-n recommendation. Thus, the personalized information distilled into the collaborative user prompt encodes more relevant information for generative recommendation compared to the task-specific prompt. However, combining the two types of prompts leads to better performance overall. Since both models are fine-tuned using same LLM (T5-small) and discrete prompt templates, we can attribute the gain in PeaPOD’s performance to the effectiveness of a collaborative user prompt that utilizes preferences shared between similar users to improve recommendation performance.
6.4. Effects of Model Size (RQ4)
We investigate the effects of a larger base model size on the recommendation performance. In particular, we compare PeaPOD-PMF (small) and PeaPOD-PMF (base), which are models fine-tuned on T5-small (60.5M params) and T5-base (223M params), respectively. We train both models with the same hyperparameters and user embeddings. We report our findings in Tables 6, 7, and 8.
For sequential recommendation, we find that PeaPOD-PMF (base) outperforms PeaPOD-PMF (small) on the Sports dataset, but underperforms on the Beauty and Toys datasets. For top-n recommendation, PeaPOD-PMF (base) performs worse for the Beauty and Toys datasets, but is competitive with PeaPOD-PMF (small) in Sports. In particular, PeaPOD-PMF (base) saw the lowest drop in performance on the Toys dataset. For explanation generation, PeaPOD-PMF (base) performs better for Sports and PeaPOD-PMF (small) performs better for Beauty. Interestingly, we find that PeaPOD-PMF (base) overfits on the model much earlier than its small counterpart on all three datasets.
Overall, we find that a bigger model does not always equate to a better performance. In fact, PeaPOD-PMF (small) model performs better on the majority of metrics despite having fewer parameters.  Geng et al. (2022) found a similar finding, noting that the smaller P5 variant outperformed the larger P5 variant on numerous tasks. In addition, both our and Geng et al. (2022)’s experiments found that the larger model performs significantly better on the Sports dataset, which has significantly more data and a lower sparsity compared to Beauty and Toys datasets. These findings indicate that a larger model may be needed when there is more data, whereas a smaller model might be more appropriate when training on a smaller dataset with higher sparsity.